40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Feb 19 15:45:43 2015, ericq, Update, CDS, Bad CDS behavior 
    Reply  Thu Feb 19 23:23:52 2015, Chris, Update, CDS, Bad CDS behavior 
       Reply  Fri Feb 20 12:08:10 2015, ericq, Update, CDS, Bad CDS behavior 
          Reply  Fri Feb 20 12:29:01 2015, ericq, Update, CDS, All optics damped 
       Reply  Fri Feb 20 14:44:47 2015, ericq, Update, CDS, Bad CDS behavior 
Message ID: 11054     Entry time: Fri Feb 20 12:29:01 2015     In reply to: 11053
Author: ericq 
Type: Update 
Category: CDS 
Subject: All optics damped 

I noticed some diagnostic bits in the c1sus IOP complaining about user application timing and FIFO under/overflow (The second and fourth squares next to the DACs on the GDS screen.) Over in crackle-cymac land, I've seen this correlate with large excess DAC noise. After restarting all models, all but one of these is green again, and the optics are now all damped. 

It seems there were some fishy BURT restores, as I found the TT control filters had their inputs and outputs switched off. Some ASS filters were found this way too. More undesired settings may still lurk in the mists...

The interferometer is now realigned, arms locked. 

ELOG V3.1.3-